When a negatively charged particle is placed in electric field, how much it will accelerate experience force? Also find distance in x and y direction?
1
Expert's answer
2011-06-09T14:48:37-0400
The force acting on the particle,
F = E/q , is directed as the electric field. The acceleration would be a = F/m = E/qm.